Circular No. 6596

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

              URL http://cfa-www.harvard.edu/cfa/ps/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVAE The Supernova Cosmology Project [cf. IAUC 6540, plus N. Walton

and P. Ruiz-Lapuente at the Isaac Newton Telescope (INT)] reports sixteen supernovae found with the Cerro Tololo (CTIO) 4-m telescope in their High Redshift Supernova Search, as given below.

SN  1997 UT  α_{2000}  δ_{2000}  R  z  type
1997ac Feb. 10 8^{h}24^{m}05.^{s}21 + 4^{o}11′22′′.6 23.1 0.32 Ia
1997ad Mar. 5 8 22 32.64 + 4 09 57.7 22.8  ?
1997ae Mar. 5 8 23 58.68 + 3 47 53.9 23.9  ?
1997af Mar. 5 8 23 52.67 + 4 08 44.0 22.3 0.58 Ia
1997ag Mar. 5 9 58 53.35 − 0 20 29.0 23.2 0.59 Ia
1997ah Mar. 5 10 00 02.92 − 0 23 30.7 23.6 0.73 Ia?
1997ai Mar. 5 10 48 57.61 + 0 31 50.0 22.3 0.45* Ia
1997aj Mar. 5 10 55 52.98 − 3 59 02.3 23.8 0.58 Ia
1997ak Mar. 5 10 56 53.15 − 4 13 50.9 24.4 0.35 Ia?
1997al Mar. 5 10 57 30.31 − 3 14 49.7 24.0 0.62 Ia
1997am Mar. 5 10 57 31.52 − 3 13 48.3 22.9 0.42 Ia
1997an Mar. 5 10 57 43.97 − 3 47 18.2 24.3 0.49 ?
1997ao Mar. 5 13 47 05.42 + 2 08 59.5 24.3  ?
1997ap Mar. 5 13 47 09.90 + 2 23 57.5 24.2 0.83* Ia
1997aq Mar. 5 13 47 09.90 + 2 27 37.2 24.3 0.51 ?
1997ar Mar. 5 13 47 31.98 + 2 26 16.0 23.0 0.36 ?

The supernovae were all discovered before or at maximum light; they were re-observed on Mar. 9–14 at CTIO, WIYN, and INT, and the supernova spectra were obtained at Mauna Kea (Keck telescope) on Mar. 13-14, except for those with no redshift indicated. The redshifts were determined from host-galaxy emission and absorption lines, except for those marked with an asterisk, which were determined solely from the spectra of the supernovae. The supernovae are all within 2′′.5 of the host-galaxy center, except for SN 1997ap (which has no detected host). Photometry is being followed in R and I for the twelve supernovae at z > 0.35, and in V, R, and I for SN 1997ac; finding charts are available from saul@LBL.gov.

Corrigendum. On IAUC 6540, line −1, for supervovae read supernovae

1997 March 20  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1997 CBAT  Daniel W. E. Green